OpenStudy (anonymous):

simplify (3+1/4)/(5/6+4). just need the formula cause i know the answer just not how to put it on paper

OpenStudy (anonymous):

first add \[3+\frac{1}{4}=\frac{13}{4}\] then add \[\frac{5}{6}+4=\frac{29}{4}\]

OpenStudy (anonymous):

then you have \[\frac{13}{4}\div\frac{29}{4}=\frac{13}{4}\times \frac{4}{29}=\frac{13}{29}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw